OpenStudy (anonymous):

2 cos^2 x - 3 cos x + 1 = 0 for 0 is less than or equal to x is less than 2pi

OpenStudy (anonymous):

the first term with the cosine... is it: \(\large 2cos^2x \) or \(\large 2cos(2x) \) i'm thinking it's the first one?

OpenStudy (anonymous):

yea the first one

OpenStudy (anonymous):

think of it this way.. let y=cosx so your equation becomes: \(\large 2y^2-3y+1=0 \) can you solve this quadratic?

OpenStudy (anonymous):

yea i can

OpenStudy (anonymous):

what is/are the solutions ? y = ???

OpenStudy (anonymous):

cos2x=2cos^2(x)-1

OpenStudy (anonymous):

2(2cos^2(x)-1)-3cosx+1=0 -> 4cos^2(x)-3cosx-1=0 -> (4cosx+1)(cosx-1)=0

OpenStudy (anonymous):

x1 = -0.5 x2 = -1

OpenStudy (anonymous):

so set 4cosx+1=0 or cos-1 =0

OpenStudy (anonymous):

hmmm.. i got the same but POSITIVE values... i'm double checking...

OpenStudy (anonymous):

yeah... they should be positive: \(\large 2y^2-3y+1=0 \) \(\large (2y-1)(y-1)=0 \) so 2y - 1= 0 gives y=1/2 y-1 = 0 gives y=1

OpenStudy (anonymous):

ok so far?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

you forgot to x2

OpenStudy (anonymous):

it is 2cos2x

OpenStudy (anonymous):

it is 2 cos^2 x - 3

OpenStudy (anonymous):

oh i thought it is cos2x

OpenStudy (anonymous):

since y=cosx, we have these two equations: \( \large cosx=\frac{1}{2}\) and \(\large cosx=-1 \) can you solve these?

OpenStudy (anonymous):

oops... that second one should be POSITIVE one...

OpenStudy (anonymous):

do you use the unit circle?

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

this is probably a stupid question, but i'm a bit confused. you said let y = cosx so what happens to the first term 2cos2x? like what about the 2 infront of the x?

OpenStudy (anonymous):

@joshi its 2 cos^2 x -3 i made a mistake in typing it

OpenStudy (anonymous):

ok... so look at what angle gives you a cosine of 1/2.... HINT: there are two of 'em from 0 to 2pi

OpenStudy (anonymous):

ohhh ok that makes more sense haha

OpenStudy (anonymous):

yeah... i got a clarification from the asker with my first question... :)

OpenStudy (anonymous):

60 degrees?

OpenStudy (anonymous):

 yep... 60 degrees is one of 'em....